OpenStudy (anonymous):

A railroad tunnel is shaped like a semiellipse as shown below. A semiellipse is shown on the coordinate plane with vertices on the x axis and one point of intersection with the positive y axis. The height of the tunnel at the center is 54 ft and the vertical clearance must be 18 ft at a point 8 ft from the center. Find an equation for the ellipse. I will post a pic in a second.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@campbell_st

OpenStudy (danjs):

hey gave 2 points, and it is centered at the origin. x^2 / a^2 + y^2 / b^2 = 1 use the two points to find a and b....

OpenStudy (anonymous):

plug in values for x and y?

OpenStudy (danjs):

yeah.. as a check , one of the points is already on the y axis

OpenStudy (danjs):

for the b value

OpenStudy (anonymous):

(0,54) is good right?

OpenStudy (danjs):

yes, and (8,18)

OpenStudy (anonymous):

0^2/a^2 so that cancels out and I have: 54^2/b^2 = 1

OpenStudy (danjs):

yeah it crosses the y axis at 54, so that checks out for b right

OpenStudy (anonymous):

2916/b^2 = 1 how do I simplify?

OpenStudy (danjs):

it will cross the y axis at + and - b

OpenStudy (danjs):

multiply by b^2, then take square root

OpenStudy (danjs):

should just say 54^2 = b^2 , so b is 54

OpenStudy (anonymous):

oh lol, which beans I get +-54?

OpenStudy (anonymous):

ok and then I plug that back in with the point 8,18 right?

OpenStudy (danjs):

yep, they really gave you that value, but it checks correct

OpenStudy (danjs):

need to find + and - a, where it crosses x axis

OpenStudy (danjs):

use the other point, and the b value, solve a

OpenStudy (anonymous):

8^2/a^2 + 18^2/54^2 = 1 64/a^2 + 324/2916= 1 how do I simplify

OpenStudy (anonymous):

I got a = 6 * sqrt(2) Is that right?

OpenStudy (anonymous):

so I get: x^2/72 + y^2/2916 = 1 for the final equation is that right?

OpenStudy (danjs):

when you have \[\frac{ 18^2 }{ 54^2 } = [\frac{ 18 }{ 54 }]^2\] reduce the inner fraction first, then square it

OpenStudy (danjs):

[1/3]^2 = 1/9

OpenStudy (anonymous):

is my final equation correct?

OpenStudy (anonymous):

x^2/72 + y^2/2916 = 1?

OpenStudy (danjs):

yes that is right for a, but remember it is + and -

OpenStudy (anonymous):

so how do I change the final equation?

OpenStudy (danjs):

for the standard form, you want to leave those denominators as a^2 and b^2, not expanded out

OpenStudy (danjs):

\[\frac{ x^2 }{ [6\sqrt{2}]^2 }+\frac{ y^2 }{ 54^2 } = 1\]

OpenStudy (anonymous):

ok thanks

OpenStudy (danjs):

welcome